# Economic Complexity Alignment and Sustainable Development

### v1

@@ 3.1 Data
Consistent with prior literature [Hartmann et al., 2017, Pinheiro et al., 2022, Stojkoski et al., 2023, Jun et al., 2020, Hartmann et al., 2021], we exclude small economies by retaining only countries with total exports exceeding 11 billion USD, and a population larger than 11 million in 2011. Additionally, we remove smaller products with global exports lower than $500$ million USD in 2011. This minimises distortions arising from the volatile specialisation patterns of small countries or products [Stojkoski et al., 2023], and results in a final sample of 138 countries and 1029 products from 2011 to 2023.
@@ 3.4 Economic Complexity
[Hidalgo, 2021, Mealy et al., 2019, Caldarelli et al., 2012, Cristelli et al., 2013]. The eigenvectors of this system associated with the second largest eigenvalues capture the largest amount of variance in the system [Hausmann and Hidalgo, 2013], and are therefore used to quantify the complexity of countries and products [^2] The eigenvectors with the second largest eigenvalues are used since $\tilde{M}_{cc^{\prime}}]$ and $\tilde{M}_{pp^{\prime}}]$ are row-stochastic matrices with a first eigenvector of 11s, hence carrying trivial information [Hidalgo, 2021, Sciarra et al., 2020, Caldarelli et al., 2012].. Because of the relative nature of the complexity metrics, these eigenvectors ($\vec{K}$ for $\tilde{M}_{cc^{\prime}}]$ and $\vec{L}$ for $\tilde{M}_{pp^{\prime}}]$) are standardised using a $Z$-transformation, resulting in the following expressions for the economic complexity index and product complexity index:
@@ 3.5 Product-Level Sustainability Measures
where $T_{c}$ is the performance score of country $c$, $s_{cp}=X_{cp}/\sum_{p^{\prime}}X_{cp^{\prime}}$ represents the share of product $p$ in the export basket of country $c$, and $M_{cp}$ denotes the elements of the binarised specialisation matrix, which ‘activates’ the export shares to only consider countries that have an $\mathit{RCA}_{cp}\geq 1$ [Hartmann et al., 2017]. $T_{p}$ is therefore an estimate of the expected value of the performance measure $T_{c}$ among heavy exporters of pp.
@@ 3.7 Product Entry Dynamics
We define an entry or successful diversification as a country developing a sustained $\mathit{RCA}_{cp}\geq 1$ in a product pp. To mitigate the risk of false positives caused by temporary fluctuations, we use both a backward and forward condition [Pinheiro et al., 2018]. The backward condition ($B_{cpt}$) mandates that RCA was smaller than 11 for $m=3$ consecutive years before the entry, while the forward condition ($F_{cpt}$) requires the country to sustain an RCA $\geq 1$ for the 33 subsequent years. These conditions make the entry measure more stable, enhancing the robustness of the results. Missing values of $M_{cpt}$ are excluded. Formally, this binary outcome is defined as
@@
where $\Lambda(x)=1/(1+\exp(x))$ is the logistic cumulative distribution function which converts the linear predictor into a probability bounded between 0 and 11. Next, $\omega_{cpt-1}$ denotes the relatedness density between country $c$ and product $p$ at time t1t-1, which accounts for the capability structure present within $c$ in the previous period. The term $\log(\mathit{RCA}_{cpt-1})$ captures path dependence in comparative advantage, recognising that RCA values below unity might still signal a trajectory toward entry. $\mathit{PCI}_{pt-1}$ is the lagged product complexity index, $\mathit{PSPI}_{pt-1}^{\perp}$ is the lagged product social progress index, and $\mathit{PEPI}_{pt-1}^{\perp}$ is the lagged product environmental performance index, both orthogonalised with respect to $\mathit{PCI}_{pt-1}$.
@@ 4.2 Diversification Bias Towards Complexity
where $\mathbb{P}=\mathbb{P}(\Delta M_{cpt}=1)$, and which corresponds to a threshold of $\omega^{\star}_{\mathit{PCI}}=-\delta_{3}/\delta_{6}$. Specifically, the estimates suggest that a higher $\mathit{PCI}_{cpt-1}$ only raises the odds of entry once relatedness density is greater than $(-0.298/1.062)\approx 0.281$. This is a relatively high barrier, considering that there were only $45$ countries in 2021 that had at least 11 unattained product with a density above this threshold.
In addition, while there is no significant direct effect of environmental performance when stripped of complexity, the coefficient of the interaction term with relatedness is significantly larger than zero. The relatedness threshold at which a 11 standard deviation increase in $\mathit{PEPI}_{pt-1}^{\perp}$ raises the odds of entry is $(-0.072/0.429)\approx 0.168$. Notably, this threshold is considerably lower than the relatedness threshold for $\mathit{PCI}_{cpt-1}$, which means that countries are more likely to diversify into more environmentally sustainable products at much lower capability requirements. Besides this, $\mathit{PSPI}_{pt-1}^{\perp}$ has neither a direct nor interaction effect that reaches significance, suggesting that variation in a product’s social progress does not drive countries’ entry decisions once product complexity is taken into account.
@@ Appendix A Variance Decomposition of Product Performance Measures Across Categories
Table 1 shows the decomposition of within versus between category variability of the three product performance measures $T_{p}$ in 2022 using a one-way analysis of variance (ANOVA). $S$ represents the sum of squared errors, and MS is the mean squared error. $\eta^{2}=\mathit{SS}_{\mathrm{between}}/\mathit{SS}_{\mathrm{total}}$ represents the share of total variance explained by category, and $\omega^{2}=\frac{\mathit{SS}_{\mathrm{between}}-\mathit{df}_{\mathrm{between}}\cdot\mathit{MS}_{\mathrm{error}}}{\mathit{SS}_{\mathrm{total}}+\mathit{MS}_{\mathrm{error}}}$ is a bias-corrected estimate of the population share of variance explained by category.

### v2

@@ 3.1 Data
Consistent with prior literature [Hartmann et al., 2017, Pinheiro et al., 2022, Stojkoski et al., 2023, Jun et al., 2020, Hartmann et al., 2021], we exclude small economies by retaining only countries with total exports exceeding $1$ billion USD, and a population larger than $1$ million in 2011. Additionally, we remove smaller products with global exports lower than $500$ million USD in 2011. This minimises distortions arising from the volatile specialisation patterns of small countries or products [Stojkoski et al., 2023], and results in a final sample of 138 countries and 1029 products from 2011 to 2023.
@@ 3.4 Economic Complexity
[Hidalgo, 2021, Mealy et al., 2019, Caldarelli et al., 2012, Cristelli et al., 2013]. The eigenvectors of this system associated with the second largest eigenvalues capture the largest amount of variance in the system [Hausmann and Hidalgo, 2013], and are therefore used to quantify the complexity of countries and products [^2] The eigenvectors with the second largest eigenvalues are used since $\tilde{M}_{cc^{\prime}}]$ and $\tilde{M}_{pp^{\prime}}]$ are row-stochastic matrices with a first eigenvector of $1$s, hence carrying trivial information [Hidalgo, 2021, Sciarra et al., 2020, Caldarelli et al., 2012].. Because of the relative nature of the complexity metrics, these eigenvectors ($\vec{K}$ for $\tilde{M}_{cc^{\prime}}]$ and $\vec{L}$ for $\tilde{M}_{pp^{\prime}}]$) are standardised using a $Z$-transformation, resulting in the following expressions for the economic complexity index and product complexity index:
@@ 3.5 Product-Level Sustainability Measures
where $T_{c}$ is the performance score of country $c$, $s_{cp}=X_{cp}/\sum_{p^{\prime}}X_{cp^{\prime}}$ represents the share of product $p$ in the export basket of country $c$, and $M_{cp}$ denotes the elements of the binarised specialisation matrix, which ‘activates’ the export shares to only consider countries that have an $\mathit{RCA}_{cp}\geq 1$ [Hartmann et al., 2017]. $T_{p}$ is therefore an estimate of the expected value of the performance measure $T_{c}$ among heavy exporters of $p$.
@@ 3.7 Product Entry Dynamics
We define an entry or successful diversification as a country developing a sustained $\mathit{RCA}_{cp}\geq 1$ in a product $p$. To mitigate the risk of false positives caused by temporary fluctuations, we use both a backward and forward condition [Pinheiro et al., 2018]. The backward condition ($B_{cpt}$) mandates that RCA was smaller than $1$ for $m=3$ consecutive years before the entry, while the forward condition ($F_{cpt}$) requires the country to sustain an RCA $\geq 1$ for the $3$ subsequent years. These conditions make the entry measure more stable, enhancing the robustness of the results. Missing values of $M_{cpt}$ are excluded. Formally, this binary outcome is defined as
@@
where $\Lambda(x)=1/(1+\exp(x))$ is the logistic cumulative distribution function which converts the linear predictor into a probability bounded between 0 and $1$. Next, $\omega_{cpt-1}$ denotes the relatedness density between country $c$ and product $p$ at time t1t-1, which accounts for the capability structure present within $c$ in the previous period. The term $\log(\mathit{RCA}_{cpt-1})$ captures path dependence in comparative advantage, recognising that RCA values below unity might still signal a trajectory toward entry. $\mathit{PCI}_{pt-1}$ is the lagged product complexity index, $\mathit{PSPI}_{pt-1}^{\perp}$ is the lagged product social progress index, and $\mathit{PEPI}_{pt-1}^{\perp}$ is the lagged product environmental performance index, both orthogonalised with respect to $\mathit{PCI}_{pt-1}$.
@@ 4.2 Diversification Bias Towards Complexity
where $\mathbb{P}=\mathbb{P}(\Delta M_{cpt}=1)$, and which corresponds to a threshold of $\omega^{\star}_{\mathit{PCI}}=-\delta_{3}/\delta_{6}$. Specifically, the estimates suggest that a higher $\mathit{PCI}_{cpt-1}$ only raises the odds of entry once relatedness density is greater than $(-0.298/1.062)\approx 0.281$. This is a relatively high barrier, considering that there were only $45$ countries in 2021 that had at least $1$ unattained product with a density above this threshold.
In addition, while there is no significant direct effect of environmental performance when stripped of complexity, the coefficient of the interaction term with relatedness is significantly larger than zero. The relatedness threshold at which a $1$ standard deviation increase in $\mathit{PEPI}_{pt-1}^{\perp}$ raises the odds of entry is $(-0.072/0.429)\approx 0.168$. Notably, this threshold is considerably lower than the relatedness threshold for $\mathit{PCI}_{cpt-1}$, which means that countries are more likely to diversify into more environmentally sustainable products at much lower capability requirements. Besides this, $\mathit{PSPI}_{pt-1}^{\perp}$ has neither a direct nor interaction effect that reaches significance, suggesting that variation in a product’s social progress does not drive countries’ entry decisions once product complexity is taken into account.
@@ Appendix A Variance Decomposition of Product Performance Measures Across Categories
Table 1 shows the decomposition of within versus between category variability of the three product performance measures $T_{p}$ in 2022 using a one-way analysis of variance (ANOVA). SS represents the sum of squared errors, and MS is the mean squared error. $\eta^{2}=\mathit{SS}_{\mathrm{between}}/\mathit{SS}_{\mathrm{total}}$ represents the share of total variance explained by category, and $\omega^{2}=\frac{\mathit{SS}_{\mathrm{between}}-\mathit{df}_{\mathrm{between}}\cdot\mathit{MS}_{\mathrm{error}}}{\mathit{SS}_{\mathrm{total}}+\mathit{MS}_{\mathrm{error}}}$ is a bias-corrected estimate of the population share of variance explained by category.